7. AUFGABE DER WOCHE

Für welche y ∈ ℝ ist die Gleichung

x^{x^{x^{x^{.^{.}}}}} = y

lösbar?

(z.B. gibt es die Lösung x = \sqrt{2} für y = 2, aber keine Lösung für y = 4.)

http://www.math.uni-bielefeld.de/∼florian/adw/.